# M. C. Rodriguez

### v5

@@ 1 Introduction
$H_{1}$ $=$ $\left(\begin{array}[]{c}h^{0}_{1}\\ h^{-}_{1}\end{array}\right)\sim({\bf 1},{\bf\bar{2}},-1),\,\ \langle H_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}v_{1}\\ 0\end{array}\right),$ (67)
$H_{2}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ h^{0}_{2}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\,\ \langle H_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v_{2}\end{array}\right).$ (72)
@@
$\Delta_{1}$$\equiv$ $\left[\left(i\sigma_{2}\right){\bf T}_{1}\right]=\left(\begin{array}[]{cc}\Delta^{0}_{1}&\frac{h^{+}_{1}}{\sqrt{2}}\\ \frac{h^{+}_{1}}{\sqrt{2}}&-H^{++}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 3},+2\right),\,\ \langle\Delta_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}V_{\Delta_{1}}&0\\ 0&0\end{array}\right),$ (77)
$\Delta_{2}$$\equiv$ $\left[\left(i\sigma_{2}\right){\bf T}_{2}\right]=\left(\begin{array}[]{cc}\Delta^{--}_{2}&\frac{h^{-}_{2}}{\sqrt{2}}\\ \frac{h^{-}_{2}}{\sqrt{2}}&-\Delta^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf\bar{3}},-2\right),\,\ \langle\Delta_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ 0&V_{\Delta_{2}}\end{array}\right).$ (82)
@@
Our numerical results are shown in Figs.(8,9). The usual scalar sector of this model is presented in [51, 54, 71, 55].
Figure 8: We show the prediction about the behaviour for the gauge boson mass $W^{\pm}$ in SUSYGR, see our Eq.(85), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ shown in the box in several colors and $v=246$ GeV. The straight lines in brown, orange and cyan are the values shown in our Eqs.(LABEL:wsmval,23,34), respectively.
Figure 9: We show the prediction about the behaviour for the $\rho$-parameter in SUSYGR, see our Eq.(85), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ shown in the box in several colors and $v=246$ GeV. Where we use the following values $\rho^{cen}=0.9998$, $\rho^{inf}=\rho^{cen}-0.0008$ and $\rho^{SM}=1$ as defined in our Eq.(32).
@@
    It is the simplest model that includes bileptons of both types: scalars and vectors ones, see our Eq.(135);
@@
    The models have a scalar sector similar to the two Higgs doublets Model, see Sec.(2) or our Eqs.(93,99,105,111), which allow to predict the quantization of the electric charge [78];
@@
A few years ago, it was made a preliminary study about the Majoron in the $m331$, in this article they do not used interactions that break lepton number conservation [45]. Therefore, they do not remove the Majoron, $M^{0}$. However in the $m331$ the Majoron, $M^{0}$, does not couple, at the tree level, neither to charged leptons nor to the quarks, then we avoid the so dangeours decay $Z^{0}\rightarrow R_{\Delta}+M^{0}$ [45, 86].
The minimal supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, was built in [87, 88, 89, 90, 91]. An interesting aspect of the model, is that the mass of the lighest Higgs is around $125.5\,\ GeV$ [89, 91]. In this preliminar analysis, the hypothesis was used that only the scalars $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now we will allow all the usual scalars of the MSUSY331 to have VEV, it means we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [48].
The goal of this work is to show that using a mechanism similar to scheme of Ma, see our Eq.(51), in MSUSY 331, we were able to remove the Majoron from the model in a similar mechanism as presented in SUSYGR. We get this nice solution, because the neutral fields belong to the sextet, ($\sigma^{\prime 0}_{1}$, $\sigma^{\prime 0}_{2}$), and to the anti-sextet, ($\sigma^{0}_{1}$), representation of the group $SU(3)_{L}$, are not necessary to generate mass for any charged fermion in the model, so their vaccum expectation values can be a few GeV, but the pseudoscalar will get large mass around ${{\cal O}}(400)$ GeV.
@@
We will introduce the following chiral superfields associated with leptons and quarks $\hat{L}_{1,2,3}$, $\hat{Q}_{1,2,3}$, $\hat{u}^{c}_{1,2,3}$, $\hat{d}^{c}_{1,2,3}$, $\hat{J}^{c}$ and $\hat{j}^{c}_{1,2}$ [89, 90]. The particle content of each chiral superfield and anti-chiral supermultiplet is presented in the Tabs.(1,2), respectively.
@@
In an recente article we explained the $W$-boson mass, presented by the CDF [^21] Compare with Eq.(85)., in the MSUSY331 [48]
@@
$\rho$ $=$ $\frac{1+R}{1+R^{\prime}},$
$R$ $=$ $\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),\,\ R^{\prime}=\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+4\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),$
$\rho$ $\simeq$ $1+R-R^{\prime}=1-2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right).$ (134)
Those result is similar with our Eq.(85).
The neutral gauge bosons are the photon
where $\theta_{W}$ is the Weinberg angle and it is defined by the following algebric relation
$\frac{M^{2}_{Z^{0}}}{M^{2}_{W^{\pm}}}\equiv\frac{1}{\cos^{2}\theta_{W}}=\frac{1}{1-\sin^{2}\theta_{W}},$ (137)
and we have also two massives neutral bosons defined as
with
$h(t)=1+4t^{2},$ (140)
for neutrinos and
for the charged leptons, where we have defined
We can use muon collider to discover the new neutral $Z^{\prime}$ boson using the reaction $\mu e\to\mu e$ it was shown at [99, 101] that $A_{RL}(\mu e)$ asymmetry is considerably enhanced, some phenomenological analyses was presented by references [89, 90, 97].
$g+d\to U^{--}+J,\,\ g+u\to U^{--}+j_{\alpha},$ (144)
its signature is $llXX$ and it can be descted at LHC if they really exist in nature [97, 105].
@@
We want to emphasize, the gauge bosons $V^{\pm}$ and $U^{\pm\pm}$ are bileptons, see our Eq.(135). The term $(\hat{L}\hat{S}\hat{L})$ will introduce a term similar to Eq.(LABEL:yukawatermogelmini) then some of our scalars are also bilepton. The terms proportional for $f_{2},f_{4},f^{\prime}_{2}$ and $f^{\prime}_{4}$ [^24] They are similar to the coefficients $\lambda_{1}$ and $\lambda_{2}$ in the SUSYGR, see our Eq.(LABEL:suppotSUSYGR). violate the conservation of the quantum number
${\cal F}=B+L,$ (148)
where $B$ is the baryon number while $L$ is the total lepton number  [^25] See our Eqs.(26,51) [89, 90, 91, 102]. However, if we assume the global $U(1)_{\cal F}$ symmetry, it allows us to introduce the $R$-conserving symmetry, in a similar way as we define $R$-parity in MSSM [68], defined as [97]
$R=(-1)^{3{\cal F}+2S}.$ (149)
The ${\cal F}$ number attribution is
@@
The soft terms to break SUSY [^26] We are not considerating the Gaugino Mass terms neither sleptons and squarks, about those soft terms you can see them in [68]. is written as
@@
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=11.19\,\ {\rm GeV}$ and the parameters, defined by Eq.(143), for this situation we got the values shown in our Tab.(4)
    $v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
    1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
    1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
    2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
    2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
    3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
    3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
    4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
    4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
    5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
@@
    In this case those VEV have to satisfy the following constraint [48]
@@
    some possible values is shown in our Tab.(5). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89\,\ GeV$, $v_{\sigma_{2}}=10\,\ GeV$, $v_{\sigma^{\prime}_{1}}=0.5\,\ GeV$ and $v_{\sigma^{\prime}_{2}}=8.77\,\ GeV$ and using the parameters defined in our Eq.(143) we obtain the values presented in our Tab.(6).
@@
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
@@ 3 Case 1
[…]
we also use the constraint [^27] To recover the VEV of SM given in our Eqs.(16,134).
@@
Figure 10: We show the prediction about the behaviour of the masses of lightest scalars, $M_{h^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ shown in the box in several colors in case 1. $M^{exp}_{h^{0}}$ (straight line Brown) is the central experimental values given at our Eq.(20), $M^{sup}_{h^{0}}=M^{exp}_{h^{0}}+0.11$ (straight line gray) while $M^{inf}_{h^{0}}=M^{exp}_{h^{0}}-0.11$ (straight line yellow).
Figure 11: We show the prediction about the behaviour of the masses of the usual heavy scalar, $M_{H^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ shown in the box in several colors in case 1. $M^{exp}_{H^{0}}$ is the lower experimental values, given in our Eq.(54), (straight line Brown).
@@
and we have verified analytically, we have two Goldstone bosons ($G^{0}_{Z^{0}},G^{0}_{Z^{\prime 0}}$). Those Goldstones bosons will give masses for the neutral gauge boson $Z^{0}$ and $Z^{\prime 0}$, respectively. The other six physical pseudoscalars have the following masses in (GeV) when we use the VEV defined in our Eq.(159), they are the following values
@@
Figure 12: We show the prediction about the behaviour of the masses of the lightest pseudoscalar, $M_{A^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ shown in the box in several colors in case 1. $M^{exp}_{A^{0}}$ is the lower experimental values, given in our Eq.(56), (straight line Brown).
@@
The possible values of the masses for $h^{0},H^{0}$ and pseudoscalar $A^{0}$ in terms of $v_{\chi}$ are shown in our Figs.(10,11,12), respectivelly. To obtain upper limits on the masses of $H^{0}$ and $A^{0}$ we draw Fig.(13), where we show $M_{H^{0}}$ and $M_{A^{0}}$ for $v_{\chi}>20$ TeV. From these figures, we can say that
@@
Therefore, our results for the masses of neutral scalars are within current experimental limits shown in our Eqs.(20,54,56).
In the Single charged Scalars we have two Goldstone bosons $G^{\pm}_{W^{\pm}}$ and $G^{\pm}_{V^{\pm}}$. They will give masses for the gauge bosons $W^{\pm}$ and $V^{\pm}$, respectively. And we also have nine massive states, in GeV, in two unrelated bases. The first base is defined as
@@
for this new basis the eigenvalues in (GeV) have the following values for the VEV given in our Eq.(159)
@@
In the Double charged Scalars we have one Goldstone bosons, $G^{\pm\pm}_{U^{\pm\pm}}$, to give mass for the gauge boson $U^{\pm\pm}$ and seven massive states, and their masses in (GeV) for the VEV given in our Eq.(159) are
@@
Figure 15: We show the prediction about the behaviour of the masses of $M_{h^{\pm\pm}}$ in terms of $v_{\chi}$ for some values of $v_{\chi^{\prime}}$ shown in the box in several colors.

### v6

@@ 1 Introduction
There is also, the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [76, 80, 78]. The field content of the MRSSM is enlarge compared to the MSSM as we shown in our Tab.(1)and the triplet $\hat{T}$ is defined as [80, 78]
$\hat{T}$ $=$ $\left(\begin{array}[]{cc}\frac{\hat{T^{0}}}{\sqrt{2}}&\hat{T}^{+}\\ \hat{T}^{-}&\frac{\hat{T^{0}}}{\sqrt{2}}\end{array}\right).$ (79)
SuperfieldField $(SU(3)_{C},SU(2)_{L})_{U(1)_{Y}}$ $U(1)_{R}$
$\hat{Q}_{L}$ $({\bf 3},{\bf 2})_{1/6}$ 1
$\hat{U}_{R}$ $({\bf\bar{3}},{\bf 1})_{-2/3}$ 1
$\hat{D}_{R}$ $({\bf\bar{3}},{\bf 1})_{1/3}$ 1
$\hat{L}_{L}$ $({\bf 1},{\bf 2})_{-1/2}$ 1
$\hat{E}_{R}$ $({\bf 1},{\bf 1})_{1}$ 1
$\hat{S}$ $({\bf 1},{\bf 1})_{0}$ 0
$\hat{T}$ $({\bf 1},{\bf 3})_{0}$ 0
$\hat{{\cal O}}$ $({\bf 8},{\bf 1})_{0}$ 0
$\hat{H}_{1}$ $({\bf 1},{\bf 2})_{1/2}$ 0
$\hat{H}_{2}$ $({\bf 1},{\bf 2})_{-1/2}$ 0
$\hat{T}$ $({\bf 1},{\bf 3})_{0}$ 0
$\hat{S}$ $({\bf 1},{\bf 1})_{0}$ 0
$\hat{R}_{1}$ $({\bf 1},{\bf 2})_{-1/2}$ 2
$\hat{R}_{2}$ $({\bf 1},{\bf 2})_{+1/2}$ 2
Table 1: Matter and $R$-charges in the $R$-symmetric supersymmetric model, MRSSM.
The superpotential of this model is [79]
$W=\left(\mu_{1}+\lambda_{1}\hat{S}\right)\hat{H}_{1}\hat{R}_{1}+\left(\mu_{2}+\lambda_{2}\hat{S}\right)\hat{H}_{2}\hat{R}_{2}+\Lambda_{1}\hat{R}_{1}\hat{T}\hat{H}_{1}+\Lambda_{2}\hat{R}_{2}\hat{T}\hat{H}_{2}.$ (80)
We can generate Majorana mass for neutrinos through the operator
$H_{2}H_{2}L_{iL}L_{jL}.$ (81)
The mass of $W$ is given by [80, 78]
$M^{2}_{W}=\frac{g^{2}}{4}\left(v^{2}+4v^{2}_{T}\right)=\frac{g^{2}v^{2}}{4}\left(1+R\right),\,\ R=\left(\frac{2v_{T}}{v}\right)^{2},$ (82)
We can explain the new measurement for the $W$-boson mass, given at Eq.(51), if $|v_{T}|\leq 4$ GeV [80, 78] and see our Fig.(8). The mass expression for $Z$ gauge boson has the same expression as presented in our Eq.(LABEL:wsmval). The $\rho$-parameter is [80, 78]
$\rho=\left(1+R\right),\,\ R=\left(\frac{2v_{T}}{v}\right)^{2},$ (83)
and see our Eq.(9). The mass spectrum in Higgs sector was presented [80, 79, 80, 81].
Figure 8: We show the prediction about the behaviour for the gauge boson mass $W^{\pm}$ in MRSSM, see our Eq.(82), as a function of the Triplet VEV $v_{T}$. The straight lines in blue, red and green are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
Figure 9: We show the prediction about the behaviour for the $\rho$-parameter in MRSSM, see our Eq.(83), as a function of the Triplet VEV $v_{T}$. Where we use the following values $\rho^{cen}=0.9998$, $\rho^{lim}_{sup}=\rho^{cen}+0.0008$ and $\rho^{SM}=1$ as defined in our Eq.(49).
@@
$H_{1}$ $=$ $\left(\begin{array}[]{c}h^{0}_{1}\\ h^{-}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right),\,\ \langle H_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}v_{1}\\ 0\end{array}\right),$ (88)
$H_{2}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ h^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right),\,\ \langle H_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v_{2}\end{array}\right).$ (93)
@@
$\Delta_{1}$ $=$ $\left(\begin{array}[]{cc}\Delta^{0}_{1}&-\frac{h^{+}_{1}}{\sqrt{2}}\\ -\frac{h^{+}_{1}}{\sqrt{2}}&-H^{++}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 3},+1\right),\,\ \langle\Delta_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}V_{\Delta_{1}}&0\\ 0&0\end{array}\right),$ (98)
$\Delta_{2}$ $=$ $\left(\begin{array}[]{cc}\Delta^{--}_{2}&-\frac{h^{-}_{2}}{\sqrt{2}}\\ -\frac{h^{-}_{2}}{\sqrt{2}}&-\Delta^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf\bar{3}},-1\right),\,\ \langle\Delta_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ 0&V_{\Delta_{2}}\end{array}\right).$ (103)
@@
Our numerical results are show in Figs.(10,11). The usual scalar sector of this model is presented in [57, 60, 79, 83]. The resolution of these two problems at the same time in this model, we believe, should be better studied later in this SUSYGR.
Figure 10: We show the prediction about the behaviour for the gauge boson mass $W^{\pm}$ in SUSYGR, see our Eq.(106), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ show in the box in several colors and $v=246$ GeV. The straight lines in brown, orange and cyan are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
Figure 11: We show the prediction about the behaviour for the $\rho$-parameter in SUSYGR, see our Eq.(106), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ show in the box in several colors and $v=246$ GeV. Where we use the following values $\rho^{cen}=0.9998$, $\rho^{inf}=\rho^{cen}-0.0008$ and $\rho^{SM}=1$ as defined in our Eq.(49).
@@
    It is the simplest model that includes bileptons of both types: scalars and vectors ones, see our Eq.(LABEL:lq);
@@
    The models have a scalar sector similar to the two Higgs doublets Model, see Sec.(2) or our Eqs.(114,120,126,132), which allow to predict the quantization of the electric charge [90];
@@
A few years ago, it was made a preliminary study about the Majoron in the $m331$, in this article they do not used interactions that break lepton number conservation [51]. Therefore, they do not remove the Majoron, $M^{0}$. However in the $m331$ the Majoron, $M^{0}$, does not couple, at the tree level, neither to charged leptons nor to the quarks, then we avoid the so dangerous decay $Z^{0}\rightarrow R_{\Delta}+M^{0}$ [51, 98].
The minimal supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, was built in [99, 100, 101, 102, 103]. An interesting aspect of the model, is that the mass of the lightest Higgs is around $125.5$ GeV [101, 103]. In this preliminary analysis, the hypothesis was used that only the scalars $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now we will allow all the usual scalars of the MSUSY331 to have VEV, it means we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [54].
The goal of this work is to show that using a mechanism similar to scheme of Ma, see our Eq.(68), in MSUSY 331, we were able to remove the Majoron from the model in a similar mechanism as presented in SUSYGR. We get this nice solution, because the neutral fields belong to the sextet, ($\sigma^{\prime 0}_{1}$, $\sigma^{\prime 0}_{2}$), and to the anti-sextet, ($\sigma^{0}_{1}$), representation of the group $SU(3)_{L}$, are not necessary to generate mass for any charged fermion in the model, so their vacuum expectation values can be a few GeV, but the pseudoscalar will get large mass around ${{\cal O}}(400)$ GeV and in this way removing the non-massive Majoron from MSUSY331.
@@
We will introduce the following chiral superfields associated with leptons and quarks $\hat{L}_{1,2,3}$, $\hat{Q}_{1,2,3}$, $\hat{u}^{c}_{1,2,3}$, $\hat{d}^{c}_{1,2,3}$, $\hat{J}^{c}$ and $\hat{j}^{c}_{1,2}$ [101, 102]. The particle content of each chiral superfield and anti-chiral supermultiplet is presented in the Tabs.(2,3), respectively.
@@
The neutral gauge bosons are the photon
where $\theta_{W}$ is the Weinberg angel and we have also two massives neutral bosons defined as
In an recent article we explained the $W$-boson mass, presented by the CDF [^17] Compare with Eq.(106)., in the MSUSY331 [54]
@@
$\rho$ $=$ $\frac{1+R}{1+R^{\prime}}\simeq 1+R-R^{\prime}=1-2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),$
$R$ $=$ $\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),\,\ R^{\prime}=\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+4\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right).$ (157)
Those result is similar with our Eq.(106). We present some phenomenological results in A.
@@
We want to emphasize, the gauge bosons $V^{\pm}$ and $U^{\pm\pm}$ are bileptons, see our Eq.(LABEL:lq). The term $(\hat{L}\hat{S}\hat{L})$ will introduce a term similar to Eq.(LABEL:yukawatermogelmini) then some of our scalars are also bilepton. The terms proportional for $f_{2},f_{4},f^{\prime}_{2}$ and $f^{\prime}_{4}$ [^18] They are similar to the coefficients $\lambda_{1}$ and $\lambda_{2}$ in the SUSYGR, see our Eq.(LABEL:suppotSUSYGR). violate the conservation of the quantum number
${\cal F}\equiv B+L,$ (161)
where $B$ is the baryon number while $L$ is the total lepton number  [^19] See our Eqs.(43,68) [101, 102, 103, 114]. However, if we assume the global $U(1)_{\cal F}$ symmetry, it allows us to introduce the $R$-conserving symmetry, in a similar way as we define $R$-parity in MSSM [74], defined as [109]
$R\equiv(-1)^{2S+3{\cal F}}.$ (162)
Where $S$ is the spin for a given state. The ${\cal F}$ number attribution is
@@
The soft terms to break SUSY [^20] We are not considering the Gaugino Mass terms neither sleptons and squarks, these soft terms are defined in [74]. is written as
@@
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10$ GeV and $v_{\sigma^{\prime}_{2}}=11.19$ GeV.
@@
    In this case those VEV have to satisfy the following
@@
    some possible values is shown in our Tab.(5). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89$ GeV, $v_{\sigma_{2}}=10$ GeV, $v_{\sigma^{\prime}_{1}}=0.5$ GeV and $v_{\sigma^{\prime}_{2}}=8.77$ GeV.
We show some details of our scalar potential in Apps.(B,C).
@@ 3 Case 1
$-k_{1}=k^{\prime}_{1}=10,\;k_{3}=k^{\prime}_{3}=-100,\;-\mu_{\eta}=\mu_{\rho}=-\mu_{s}=\mu_{\chi}=1000\quad GeV,$
we also use the constraint [^21] To recover the VEV of SM given in our Eqs.(24,157).
@@
Figure 12: We show the prediction about the behaviour of the masses of lightest scalars, $M_{h^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{h^{0}}$ (straight line Brown) is the central experimental values given at our Eq.(28), $M^{sup}_{h^{0}}=M^{exp}_{h^{0}}+0.11$ (straight line gray) while $M^{inf}_{h^{0}}=M^{exp}_{h^{0}}-0.11$ (straight line yellow).
Figure 13: We show the prediction about the behaviour of the masses of the usual heavy scalar, $M_{H^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{H^{0}}$ is the lower experimental values, given in our Eq.(71), (straight line Brown).
@@
and we have verified analytically, we have two Goldstones bosons ($G^{0}_{Z^{0}},G^{0}_{Z^{\prime 0}}$). These Goldstones bosons will give masses for the neutral gauge boson $Z^{0}$ and $Z^{\prime 0}$, respectively. The other six physical pseudoscalars have the following masses in (GeV) when we use the VEV defined in our Eq.(172), they are the following values
@@
Figure 14: We show the prediction about the behaviour of the masses of the lightest pseudoscalar, $M_{A^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{A^{0}}$ is the lower experimental values, given in our Eq.(73), (straight line Brown).
@@
Possible values of the masses of scalars $h^{0},H^{0}$ and pseudoscalar $A^{0}$ in terms of $v_{\chi}$ are show in our Figs.(12,13,14), respectively. To obtain upper limits on the masses of $H^{0}$ and $A^{0}$ we draw Fig.(15), where we show $M_{H^{0}}$ and $M_{A^{0}}$ for $v_{\chi}>20$ TeV. From these figures we can say that
@@
Therefore, our results for the masses of neutral scalars are within current experimental limits show in our Eqs.(28,71,73).
In our opinion, it is interesting to do a phenomenological study regarding the decays of the scalars $h^{0}$, $H^{0}$ and pseudoscalar $A^{0}$ in MSUSY331. A first series studying the decay of scalars in the 331 model without SUSY was presented in [118, 119].
It would be, also, so interesting to diagonalize these matrices in details. It means, we must to study the perturbative and unitary conditions for the potential; verify the Vacuum stability and of course the minimization condition. In our opinion, those analyses must be done first in the (m331) and then make in the MSUSY331. We have previously done a similar analysis in [^22] I am still in debt with Prof. Aahron Davidson, who introduced me to this interesting model, which should be better studied. [120].
In the Single charged Scalars we have two Goldstones bosons $G^{\pm}_{W^{\pm}}$ and $G^{\pm}_{V^{\pm}}$. They will give masses for the gauge bosons $W^{\pm}$ and $V^{\pm}$, respectively. And we also have nine massive states in two unrelated bases. The first base is defined as
@@
for this new basis the eigenvalues (GeV) have the following values for the VEV given in our Eq.(172)
@@
In the Double charged Scalars we have one Goldstone boson, $G^{\pm\pm}_{U^{\pm\pm}}$, to give mass for the gauge boson $U^{\pm\pm}$ and seven massive states, and their masses in (GeV) for the VEV given in our Eq.(172) are
@@
Figure 17: We show the prediction about the behaviour of the masses of $M_{h^{\pm\pm}}$ in terms of $v_{\chi}$ for some values of $v_{\chi^{\prime}}$ show in the box in several colors.